A mixture containing 10% ethanol (E) and 90% H2O (W) by weight is fed into a distillation column at the rate of 106.26 kg/h. The distillate (top product) contains 94.01 % ethanol and the distillate is produced at a rate of one tenth of that of the feed. 

Draw and label a flow chart of the process before solving the problem. Perform material (total) and ethanol balances to calculate the mass fraction of ethanol in the bottom product to at least four decimal places. 




1
Expert's answer
2021-09-07T23:53:30-0400

Sample flow chart




mass of mixture = 106.26kg/hr

10% of the mixture is ethanol = 10.626 kg

number of moles of ethanol = "\dfrac{10626\text{g} }{46\text{g/mol}}"

= 231 moles


90% of the mixture is water = 95.634 kg

number of moles of water = "\dfrac{95634\text{g}}{18\text{g/mol}}"

= 5313 moles


Total number of moles = 231 + 5313 = 5544 moles


mole fraction of ethanol = 231/5544 = 0.0417

mole fraction of water = 5313/5544 = 0.9583